Find the equation of the line that is parallel to y = 6x – 15 and passes through the origin.
Can someone please help? Thank you

Answer by Mathtut(3670)

lines that are parallel have the same slope
which is 6 in this equation
now using the point slope formula which is y-k=m(x-h) where (h,k) is a point on the line so y-0=6(x-0)---> y=6x